# Periodic nonlinear Schrödinger equation with distributional potential and invariant measures

### v1

@@ Abstract
for any $\lambda>0$ with $c_{p}>0$ a positive constant and $p>2$. In the focusing case, one needs to introduce a cut-off $B>0$ following [6, 16] with the measure
$\nu_{B}(\mathrm{d}u)=\frac{\mathds{1}_{\|u\|_{L^{2}}\leq B}}{Z_{B}}\exp\Big{(}(1-\lambda_{1})\int_{\mathbb{T}}|u(x)|^{2}\mathrm{d}x-\lambda\int_{\mathbb{T}}|u(x)|^{p}\mathrm{d}x\Big{)}\mu(\mathrm{d}u)$ (4.22)
@@
For $p<6$ and any $B>0$, we have
$\mathds{1}_{\|u\|_{L^{2}}\leq B}\exp\Big{(}(1-\lambda_{1})\int_{\mathbb{T}}|u(x)|^{2}\mathrm{d}x-\lambda\int_{\mathbb{T}}|\varphi(x)|^{p}\mathrm{d}x\Big{)}\in L^{1}(\mu).$ (4.23)
The result still holds for $p=6$ and $B$ small enough.
@@
$\int_{\mathbb{T}}|u(x)|^{p}\mathrm{d}x\leq C\|u\|_{H^{\frac{p-2}{2p}}}^{p}$ (4.26)
@@
$\int_{\mathbb{T}}|u(x)|^{p}\mathrm{d}x\leq C\|u\|_{L^{2}}^{(1-\theta)p}\|u\|_{H^{\sigma}}^{\theta p}$ (4.27)
with $\theta\in(0,1)$ and $\sigma\theta=\frac{p-2}{2p}$. Since we need $\sigma<\frac{1}{2}$ as well as $\sigma p<2$ in order to conclude with Young inequality, this imposes $p<4$. To extend the result to $p<6$, one needs interpolation with spaces based on $L^{q}$ with $q>2$. Besov injections and interpolation give
$\int_{\mathbb{T}}|u(x)|^{p}\mathrm{d}x$ $\lesssim\|u\|_{L^{2}}^{2}\|u\|_{L^{\infty}}^{p-2}$
$\lesssim\|u\|_{L^{2}}^{2}\|u\|_{B_{q,q}^{\frac{1}{q}+\varepsilon}}^{p-2}$
$\lesssim\|u\|_{L^{2}}^{2}\|u\|_{B_{2,2}^{0}}^{(1-\theta)(p-2)}\|u\|_{B_{\infty,\infty}^{\sigma}}^{\theta(p-2)}$
@@
$\theta(p-2)>\frac{1+2\varepsilon}{2}(p-2).$ (4.29)
Since our goal is to have $\theta(p-2)<2$, this yields the condition
$\frac{1}{2}(p-2)<2\iff p<6$ (4.30)
@@
$\mathbb{E}\Big{[}\mathds{1}_{\|u\|_{L^{2}}\leq B}e^{\int_{\mathbb{T}}|u(x)|^{p}\mathrm{d}x}\Big{]}$ $\leq\mathbb{E}\Big{[}e^{CB^{2+(1-\theta)(p-2)}\|u\|_{C^{\sigma}}^{\theta(p-2)}}\Big{]}$
$\leq\mathbb{E}\Big{[}e^{C\frac{1}{q^{\prime}}B^{2q^{\prime}+(1-\theta)(p-2)q^{\prime}}+\frac{1}{q}\|u\|_{C^{\sigma}}^{\theta(p-2)q}}\Big{]}$
$\leq e^{C\frac{1}{q^{\prime}}B^{2q^{\prime}+(1-\theta)(p-2)q^{\prime}}}\mathbb{E}\Big{[}e^{\frac{C_{\varepsilon}}{q}\|u\|_{L^{2}}^{2}+\frac{\varepsilon}{q}\|u\|_{C^{\sigma^{\prime}}}^{2}}\Big{]}$
$\leq e^{C\frac{1}{q^{\prime}}B^{2q^{\prime}+(1-\theta)(p-2)q^{\prime}}+\frac{C_{\varepsilon}}{q}B^{2}}\mathbb{E}\Big{[}e^{\frac{\varepsilon}{q}\|u\|_{C^{\sigma^{\prime}}}^{2}}\Big{]}$
for $q=\frac{1}{\theta(p-2)}>2$, $q^{\prime}$ its conjugated exponent, any $\varepsilon>0$, a constant $C_{\varepsilon}>0$ large enough and $\sigma<\sigma^{\prime}<\frac{1}{2}$ together with Young inequality. Using Fernique’s theorem and that $\mu$ is a Gaussian measure on the Banach space $C^{\sigma^{\prime}}$, we get
$\mathbb{E}\Big{[}\mathds{1}_{\|u\|_{L^{2}}\leq B}e^{\int_{\mathbb{T}}|u(x)|^{p}\mathrm{d}x}\Big{]}<\infty$ (4.31)
for any $B\geq 0$ and $p<6$. The results for $p=6$ follows for $B$ small enough.
@@
Since $(\mathds{1}_{Y\leq B}e^{X})^{q}=\mathds{1}_{Y\leq B}e^{qX}$, this also implies the result for $L^{q}$ insteand of $L^{1}$ with a condition on $B$ depending on $q$ for $p=6$. An alternative definition of the measure could have been with a density depending on $V$ with respect to the case $V=0$ studied in [6, 16, 21]. Denoting as $\nu_{B}^{0}$ the case $V=0$, we have
@@
using duality, product and injection of Besov spaces from Proposition A.1. One can also prove that $Z_{V}$ is finite using the cut-off, this is also an alternative proof of Proposition 4.1. Since the optimal parameter $B$ is known in the case $V=0$ from [21], the same is true here.
Since the measure $\mu$ is supported in $C^{\frac{1}{2}-\varepsilon}(\mathbb{T})$ for any $\varepsilon>0$, so is $\nu_{B}$ with the previous proposition and Theorem 3.4 guarantees that there exist almost surely a local solution for initial data distributed as $\nu_{B}$ in the focusing case, or $\nu$ in the defocusing case. In order to prove that the measure is invariant, we will show invariance for a truncated dynamic and pass to the limit following Bourgain [6]. In particular, this allows to construct global solutions. Recall that the truncated dynamic
$i\partial_{t}u_{N}=\mathcal{H}u_{N}+\lambda\Pi_{N}(|u_{N}|^{p-2}u_{N})$ (4.36)
@@
$\Omega_{N}:=\Pi_{N}L^{2}(\mathbb{T}).$ (4.37)
We only consider the focusing case which is harder, the defocusing case following from the same type of arguments. We first prove the following result.
Proposition 4.4.
Let $\sigma\in(\sigma_{\kappa},\frac{1}{2})$ and $2\leq p\leq 6$. For any $\delta>0$, there exist a subset $\Omega_{1}$ of $\Omega_{N}$ such that $\nu(\Omega_{N}\backslash\Omega_{1})<\delta$ and for $u_{0}\in\Omega_{1}$, there exists a unique global solution $u_{N}$ to (3.38) which satisfies
$\|u_{N}(t)\|_{H^{\sigma}}\leq C\Big{(}\log\frac{1+|t|}{\delta}\Big{)}^{\sigma^{+}}$ (4.38)
for any $t\in\mathbb{R}$ and $\sigma^{+}>\sigma$ with $C>0$ a constant depending on $p$ and $\sigma$.
Proof : For $\sigma\in(\sigma_{\kappa},\frac{1}{2})$ and $A>0$, let
$B_{N,A}^{\sigma}:=\{u\in\Omega_{N}\ ;\ \|u\|_{H^{\sigma}}\leq A\}.$ (4.39)
We proved with Proposition 3.5 that for any $u_{0}\in B_{N,A}^{\sigma}$, there exists a local solution $u_{N}$ to the truncated equation up to a time of order $A^{-\frac{6(p-2)}{p-8}}$. Then equation (3.38) is a finite dimensional system with values in $\Omega_{N}$ and Liouville’s Theorem implies the invariance of the Lebesgue measure. Using the conservation of the mass and the truncated Hamiltonian
$\mathcal{E}_{N}(u)=\frac{1}{2}\langle\mathcal{H}u,u\rangle+\frac{\lambda}{p}\int_{\mathbb{T}}|\Pi_{N}u(x)|^{p}\mathrm{d}x,$ (4.40)
this implies the invariance of the truncated Gibbs measure
$\nu_{N}(\mathrm{d}u)=\frac{1}{Z_{B,N}}\mathds{1}_{\|u\|_{L^{2}}\leq B}e^{-\mathcal{E}_{N}(u)}\prod_{n=1}^{N}\mathrm{d}u_{n}.$ (4.41)
Let $\tau_{A}=C_{0}A^{-\frac{6(p-2)}{p-8}}$ be the time of existence from Proposition 3.5 and consider the map
$S_{\tau_{A}}:\left|\begin{array}[]{ccc}B_{N,A}^{\sigma}&\to&\Omega_{N}\\ u_{0}&\mapsto&u(\tau_{A})\end{array}\right.$ (4.42)
which satisfies $\|S_{\tau_{A}}(u_{0})\|_{H^{\sigma}}\leq 2\|u_{0}\|_{H^{\sigma}}$. For any $m\geq 1$, introduce the set
$\Omega_{N,A,\sigma}^{m}:=B_{N,A}^{\sigma}\cap S_{\tau_{A}}^{-1}(B_{N,A}^{\sigma})\cap\ldots\cap S_{\tau_{A}}^{-m}(B_{N,A}^{\sigma}).$ (4.43)
Since the measure $\nu_{N}$ is invariant by the map $S_{\tau_{A}}$, we have
$\nu_{N}(\Omega_{N}\backslash\Omega_{N,A,\sigma}^{m})\leq(m+1)C_{1}e^{-C_{2}A^{2}}$ (4.44)
for positive constants $C_{1},C_{2}>0$ using that $\nu_{N}$ is absolutely continuous with respect to $\mu_{N}$ which is a Gaussian measure thus has exponential tails. For initial data $u_{0}\in\Omega_{N,A,\sigma}^{m}$, there exists a solution up to time $m\tau_{A}=mC_{0}A^{-\frac{6(p-2)}{p-8}}$ which satisfies $\|u(j\tau_{A})\|_{\mathcal{H}^{\sigma}}\leq A$ for any $j\in\llbracket 0,m\rrbracket$. For any $T>0$ and $\delta\in(0,1)$, consider the set
$\Omega_{N,\sigma}^{\delta,T}:=\Omega_{N,A,\sigma}^{m}$ (4.45)
with $A=C\sqrt{\log\big{(}\frac{1+T}{\delta}\big{)}}$ and $m=\lfloor TC_{0}^{-1}A^{\frac{6(p-2)}{p-8}}\rfloor$ and $C>0$ a large enough constant such that $(m+1)C_{1}e^{-C_{2}A^{2}}\leq\delta$. Then for any $u_{0}\in\Omega_{N,\sigma}^{\delta,T}$, there exists a solution $u_{N}$ to equation (3.38) to time $T$ such that
$\|u_{N}(T)\|_{H^{\sigma}}\leq C\sqrt{\log\big{(}\frac{1+T}{\delta}\big{)}}$ (4.46)
and we have
$\nu_{N}(\Omega_{N}\backslash\Omega_{N,\sigma}^{\delta,T})\leq\delta.$ (4.47)
The proof is complete if one takes
$\Omega_{1}=\bigcap_{n\geq 1}\Omega_{N,\sigma}^{2^{-n}\delta,2^{n}}$
using an interpolation between the $H^{\frac{1}{2}-\varepsilon}$ norm and the mass conservation for $\varepsilon$ small enough since $\sigma$ is any exponent in $(\sigma_{\kappa},\frac{1}{2})$.
$\square$
Using the convergence of the truncated dynamic to the solution of equation (1.1), we get the following proposition with $\Omega=L^{2}(\mathbb{T})$.
Proposition 4.5.
Let $\sigma\in(\sigma_{\kappa},\frac{1}{2})$ and $2\leq p\leq 6$. For any $\delta>0$, there exist a subset $\Omega_{1}$ of $\Omega$ such that $\nu(\Omega\backslash\Omega_{1})<\delta$ and for $u_{0}\in\Omega_{1}$, there exists a global solution to (1.1) which satisfies
$\|u(t)\|_{H^{\sigma}}\leq C\Big{(}\log\frac{1+|t|}{\delta}\Big{)}^{\sigma_{+}}$ (4.48)
for any $t\in\mathbb{R}$ and $\sigma_{+}>\sigma$.
Proof : Let $T>0$ and $\delta>0$. First, note that for any set $S\subset\Omega$, we have
$\nu(S)\leq C_{2}\mu(S)^{\frac{1}{2}}$ (4.49)
with $C_{2}>1$ the $L^{2}(\mu)$ norm of the density of $\nu$ with respect to $\mu$ which is finite from Proposition 4.2. For any $\sigma_{\kappa}<\sigma<\sigma_{+}<\frac{1}{2}$, the previous proposition gives a set $\Omega_{1,N}\subset\Omega_{N}$ such that $\nu(\Omega_{N}\backslash\Omega_{1,N})\leq\delta$ and that equation (3.38) admits a unique global solution $u_{N}$ for any $u_{0}\in\Omega_{1,N}$ with
$\|u_{N}(t)\|_{H^{\sigma}}\leq C\Big{(}\log\frac{1+|t|}{\delta}\Big{)}^{\sigma_{+}}$ (4.50)
for any $t\in\mathbb{R}$ with a constant $C>0$. Let
$\Omega_{1,N}^{\delta}:=\{u\in H^{\sigma}\ ;\ \|u\|_{H^{\sigma}}\leq A\quad\text{and}\quad\Pi_{N}u\in\Omega_{1,N}\}$ (4.51)
with $A=C_{1}\Big{(}\log\frac{1}{\delta}\Big{)}^{\sigma_{+}}$. Recall that $B_{N,A}^{\sigma}$ denotes the unit ball of $\Pi_{N}L^{2}(\mathbb{T})$ with respect to the $H^{\sigma}$ norm and consider $\Omega_{>N}:=(\text{Id}-\Pi_{N})L^{2}(\mathbb{T})$ and
$B_{>N,A}^{\sigma}:=\{u\in(\text{Id}-\Pi_{N})L^{2}(\mathbb{T})\ ;\ \|u\|_{H^{\sigma}}\leq A\}.$ (4.52)
We have
$\nu(\Omega\backslash\Omega_{1,N}^{\delta})$ $\leq C_{2}\sqrt{\mu(\Omega\backslash\Omega_{1,N}^{\delta})}$ (4.53)
$\leq C_{2}\sqrt{\mu(\Omega\backslash[(\Omega_{1,N}\cap B_{N,A}^{\sigma})\sqcup B_{>N,A}^{\sigma}])}$ (4.54)
$\leq C_{2}\sqrt{\mu(\Omega_{N}\backslash(\Omega_{1,N}\cap B_{N,A}^{\sigma}))\mu(\Omega_{>N}\backslash B_{>N,A}^{\sigma}])}$ (4.55)
$\leq C_{2}\sqrt{(\delta+C\delta)C\delta\varepsilon(\delta)}$ (4.56)
$\leq\delta$ (4.57)
for $\delta>0$ small enough, using that $\mu$ is a product of probability measures on eigenspaces, $C_{2}\geq 1$ the $L^{2}(\mathrm{d}\mu)$ norm of the density of $\nu$, finite by Proposition 4.2 and that $\varepsilon(\delta)$ goes to 00 as $\delta$ goes to 00 since $\sigma_{+}<\frac{1}{2}$. For $u_{0}\in\Omega_{1,N}^{\delta}$, there exists a unique solution $u$ to equation (1.1) up to time $\tau_{A}=CA^{-c(p)}$ and we have
$\sup_{t\in[0,\tau_{A}]}\|u(t)-u_{N}(t)\|_{H^{\sigma^{\prime}}}\lesssim N^{\sigma^{\prime}-\sigma}(1+A^{p-2})$ (4.58)
from Proposition 3.5 for $\sigma^{\prime}>\sigma$. We are going to propagate this error up to time $T$. Assume for $t\leq t_{0}$ that we proved
$\|u(t)-u_{N}(t)\|_{H^{\sigma^{\prime}}}\leq\varepsilon<1$ (4.59)
and consider $v_{N}$ the solution to the untruncated equation (1.1) on $[t_{0},t_{0}+\tau_{A}]$ with $v_{N}(t_{0})=u_{N}(t_{0})$. We bound $u-u_{N}$ using
$\|u(t)-u_{N}(t)\|_{H^{\sigma^{\prime}}}\leq\|u(t)-v_{N}(t)\|_{H^{\sigma^{\prime}}}+\|v_{N}(t)-u_{N}(t)\|_{H^{\sigma^{\prime}}}$ (4.60)
for $t\in[t_{0},t_{0}+\tau_{A}]$. For the first term, we have
$\sup_{t\in[t_{0},t_{0}+\tau_{A}]}\|u(t)-v_{N}(t)\|_{H^{\sigma^{\prime}}}\leq 2\|u(t_{0})-v_{N}(t_{0})\|_{H^{\sigma^{\prime}}}\leq 2\varepsilon$ (4.61)
using Theorem 3.4. For the second term, this follows from Proposition 3.5 and we have
$\sup_{t\in[t_{0},t_{0}+\tau_{A}]}\|v_{N}(t)-u_{N}(t)\|_{H^{\sigma^{\prime}}}\lesssim N^{\sigma^{\prime}-\sigma}(1+\|u_{N}(t_{0})\|_{H^{\sigma}}^{p-2}).$ (4.62)
Using that $\Pi_{N}u_{0}\in\Omega_{1}$, we get $\|u_{N}(t_{0})\|_{H^{\sigma}}\leq C\big{(}\log\frac{1+T}{\delta}\big{)}^{\sigma_{+}}$ for $t_{0}\leq T$ hence
$\|u(t)-u_{N}(t)\|_{H^{\sigma^{\prime}}}\leq 2\varepsilon+N^{\sigma^{\prime}-\sigma}\Big{(}1+C^{p-2}\big{(}\log\frac{1+T}{\delta}\big{)}^{(p-2)\sigma_{+}}\Big{)}$ (4.63)
for any $t\in[t_{0},t_{0}+\tau_{A}]$. Set $t_{j}=j\tau_{A}$ and $\delta_{j}=\|u(j\tau_{A})-u_{N}(j\tau_{A})\|_{H^{\sigma^{\prime}}}$. We have $\delta_{0}\leq N^{\sigma^{\prime}-\sigma}A$ and the previous computations yield
$\delta_{j+1}\lesssim 2\delta_{j}+N^{\sigma^{\prime}-\sigma}\Big{(}1+C^{p-2}\big{(}\log\frac{1+T}{\delta}\big{)}^{(p-2)\sigma_{+}}\Big{)}$ (4.64)
for $\tau j\leq T$ hence
$\|u(T)-u_{N}(T)\|_{H^{\sigma^{\prime}}}\lesssim C^{J+1}N^{\sigma^{\prime}-\sigma}\Big{(}1+C^{p-2}\big{(}\log\frac{1+T}{\delta}\big{)}^{(p-2)\sigma_{+}}\Big{)}+C^{J}N^{\sigma^{\prime}-\sigma}A$ (4.65)
for $J=\lfloor\frac{T}{\tau_{A}}\rfloor$ and $C>0$ a positive constant. Taking $N$ large enough depending on $T$ and $\delta$ gives
$\|u(T)\|_{H^{\sigma}}\leq C\Big{(}\log\frac{1+T}{\delta}\Big{)}^{\sigma_{+}}+1$ (4.66)
hence $u$ is well-defined up to $T$. The proof is complete with
$\Omega_{1}=\bigcap_{n\geq 1}\Omega_{1,N_{n}}^{2^{-n}\delta}$
with $N_{n}$ an increasing sequence to get a global solution.
$\square$
This finally gives the following theorem with the invariance of the measure.
@@
Proof : Let $S\subset H^{\sigma}$ be a bounded measurable set with $\sigma\in(\sigma_{\kappa},\frac{1}{2})$. We have
$|\nu(\Phi_{t}S)-\nu(S)|\leq|\nu(\Phi_{t}S)-\nu_{N}(\Phi_{t}^{N}\Pi_{N}S)|+|\nu_{N}(\Phi_{t}^{N}\Pi_{N}S)-\nu_{N}(\Pi_{N}S)|+|\nu_{N}(\Pi_{N}S)-\nu(S)|$ (4.67)
for anu $t\in\mathbb{R}$ and $N\geq 1$. The second term is equal to zero by invariante of $\nu_{N}$ under the truncated flow $\Phi_{t}^{N}$. For the third term, we have
$|\nu_{N}(\Pi_{N}S)-\nu(S)|\leq|\nu_{N}(\Pi_{N}S)-\nu(\Pi_{N}S)|+|\nu(\Pi_{N}S)-\nu(S)|$ (4.68)
where as $N$ goes to infinity, the first term goes to 00 with Proposition 4.2 and the second term with dominated convergence. The same holds for the first term with
$|\nu(\Phi_{t}S)-\nu_{N}(\Phi_{t}^{N}\Pi_{N}S)|\leq|\nu(\Phi_{t}S)-\nu(\Phi_{t}^{N}\Pi_{N}S)|+|\nu(\Phi_{t}^{N}\Pi_{N}S)-\nu_{N}(\Phi_{t}^{N}\Pi_{N}S)|$ (4.69)
hence the proof is complete for bounded $S$. The general result follows from the fact that
$\bigcup_{n\geq 0}\Omega_{1,2^{-n}}$ (4.70)
is an event of full measure for $\nu$ in $L^{2}(\mathbb{T})$ with $\Omega_{1,2^{-n}}\subset H^{\sigma}$ the set given by the previous proposition with $\delta=2^{-n}$ for any $\sigma<\frac{1}{2}$.

### v2

@@ Abstract
for any $\lambda>0$ with $c_{p}>0$ a positive constant and $m>2$. In the focusing case, one needs to introduce a cut-off $B>0$ following [6, 18] with the measure
$\nu_{B}(\mathrm{d}u)=\frac{\mathds{1}_{\|u\|_{L^{2}}\leq B}}{Z_{B}}\exp\Big{(}(1-\lambda_{1})\int_{\mathbb{T}}|u(x)|^{2}\mathrm{d}x-\lambda\int_{\mathbb{T}}|u(x)|^{m}\mathrm{d}x\Big{)}\mu(\mathrm{d}u)$ (4.22)
@@
For $m<6$ and any $B>0$, we have
$\mathds{1}_{\|u\|_{L^{2}}\leq B}\exp\Big{(}(1-\lambda_{1})\int_{\mathbb{T}}|u(x)|^{2}\mathrm{d}x-\lambda\int_{\mathbb{T}}|\varphi(x)|^{m}\mathrm{d}x\Big{)}\in L^{1}(\mu).$ (4.23)
The result still holds for $m=6$ and $B$ small enough.
@@
$\int_{\mathbb{T}}|u(x)|^{m}\mathrm{d}x\leq C\|u\|_{H^{\frac{m-2}{2m}}}^{m}$ (4.26)
@@
$\int_{\mathbb{T}}|u(x)|^{m}\mathrm{d}x\leq C\|u\|_{L^{2}}^{(1-\theta)m}\|u\|_{H^{\sigma}}^{\theta m}$ (4.27)
with $\theta\in(0,1)$ and $\sigma\theta=\frac{m-2}{2m}$. Since we need $\sigma<\frac{1}{2}$ as well as $\sigma m<2$ in order to conclude with Young inequality, this imposes $m<4$. To extend the result to $m<6$, one needs interpolation with spaces based on $L^{q}$ with $q>2$. Besov injections and interpolation give
$\int_{\mathbb{T}}|u(x)|^{m}\mathrm{d}x$ $\lesssim\|u\|_{L^{2}}^{2}\|u\|_{L^{\infty}}^{m-2}$
$\lesssim\|u\|_{L^{2}}^{2}\|u\|_{B_{q,q}^{\frac{1}{q}+\varepsilon}}^{m-2}$
$\lesssim\|u\|_{L^{2}}^{2}\|u\|_{B_{2,2}^{0}}^{(1-\theta)(m-2)}\|u\|_{B_{\infty,\infty}^{\sigma}}^{\theta(m-2)}$
@@
$\theta(m-2)>\frac{1+2\varepsilon}{2}(m-2).$ (4.29)
Since our goal is to have $\theta(m-2)<2$, this yields the condition
$\frac{1}{2}(m-2)<2\iff m<6$ (4.30)
@@
$\mathbb{E}\Big{[}\mathds{1}_{\|u\|_{L^{2}}\leq B}e^{\int_{\mathbb{T}}|u(x)|^{m}\mathrm{d}x}\Big{]}$ $\leq\mathbb{E}\Big{[}e^{CB^{2+(1-\theta)(m-2)}\|u\|_{C^{\sigma}}^{\theta(m-2)}}\Big{]}$
$\leq\mathbb{E}\Big{[}e^{C\frac{1}{q^{\prime}}B^{2q^{\prime}+(1-\theta)(m-2)q^{\prime}}+\frac{1}{q}\|u\|_{C^{\sigma}}^{\theta(m-2)q}}\Big{]}$
$\leq e^{C\frac{1}{q^{\prime}}B^{2q^{\prime}+(1-\theta)(m-2)q^{\prime}}}\mathbb{E}\Big{[}e^{\frac{C_{\varepsilon}}{q}\|u\|_{L^{2}}^{2}+\frac{\varepsilon}{q}\|u\|_{C^{\sigma^{\prime}}}^{2}}\Big{]}$
$\leq e^{C\frac{1}{q^{\prime}}B^{2q^{\prime}+(1-\theta)(m-2)q^{\prime}}+\frac{C_{\varepsilon}}{q}B^{2}}\mathbb{E}\Big{[}e^{\frac{\varepsilon}{q}\|u\|_{C^{\sigma^{\prime}}}^{2}}\Big{]}$
for $q=\frac{1}{\theta(m-2)}>2$, $q^{\prime}$ its conjugated exponent, any $\varepsilon>0$, a constant $C_{\varepsilon}>0$ large enough and $\sigma<\sigma^{\prime}<\frac{1}{2}$ together with Young inequality. Using Fernique’s theorem and that $\mu$ is a Gaussian measure on the Banach space $C^{\sigma^{\prime}}$, we get
$\mathbb{E}\Big{[}\mathds{1}_{\|u\|_{L^{2}}\leq B}e^{\int_{\mathbb{T}}|u(x)|^{m}\mathrm{d}x}\Big{]}<\infty$ (4.31)
for any $B\geq 0$ and $m<6$. The results for $m=6$ follows for $B$ small enough.
@@
Since $(\mathds{1}_{Y\leq B}e^{X})^{q}=\mathds{1}_{Y\leq B}e^{qX}$, this also implies the result for $L^{q}$ insteand of $L^{1}$ with a condition on $B$ depending on $q$ in the case $m=6$. An alternative definition of the measure could have been with a density depending on $V$ with respect to the case $V=0$ studied in [6, 18, 23]. Denoting as $\nu_{B}^{0}$ the case $V=0$, we have
@@
using duality, product and injection of Besov spaces from Proposition A.1. One can also prove that $Z_{V}$ is finite using the cut-off, this is also an alternative proof of Proposition 4.1. Since the optimal parameter $B$ is known in the case $V=0$ from [23], the same is true here. This construction of the measure is less suited to the study of its truncation associated to the operator $\mathcal{H}$ and the Schrödinger equation.
Since the measure $\mu$ is supported in $C^{\frac{1}{2}-\varepsilon}$ for any $\varepsilon>0$, this also holds for $\nu_{B}$ as it is absolutely continuous with respect to $\mu$ with the previous proposition. Then Theorem 3.4 guarantees that there exists almost surely a local solution to (1.1) for initial data distributed as $\nu_{B}$ in the focusing case, or $\nu$ in the defocusing case. We now prove the the measure is invariant, thus allowing to construct global solutions as done by Bourgain [6]. However, we rely here on a different argument from Da Prato and Debussche in the parabolic case. Recall that the truncated dynamic
$i\partial_{t}u_{N}=\mathcal{H}u_{N}+\lambda\Pi_{N}(|u_{N}|^{m-2}u_{N})$ (4.36)
@@
$\Omega_{N}:=\Pi_{N}L^{2}.$ (4.37)
Using Proposition 3.5, we prove that one can pass to the limit and get global solution with invariant measure for (1.1).
@@
Proof : Let $u_{0}$ be a random initial data with law $\nu_{B}$ in the focusing case and $\nu$ in the defocusing case. Then there exists a maximal solution up to a random time $T^{*}(u_{0})\in(0,+\infty]$ with the explosion criterion
$T^{*}(u_{0})<\infty\quad\implies\quad\sup_{t<T^{*}(u_{0})}\|u(t)\|_{H^{\sigma}}=+\infty.$ (4.38)
Let $T>0$. For any $t<T\wedge T^{*}(u_{0})$, the mild formulation for equation (1.1) gives
$\sup_{t<T\wedge T^{*}(u_{0})}\|u(t)\|_{H^{\sigma}}\lesssim\|u_{0}\|_{H^{\sigma}}+\int_{0}^{T\wedge T^{*}(u_{0})}\|u(t)\|_{L^{\infty}}^{m-2}\|u(t)\|_{H^{\sigma}}\mathrm{d}t$ (4.39)
hence Fubini-Tonelli Theorem gives
$\mathbb{E}\Big{[}\sup_{t\leq T\wedge T^{*}(u_{0})}\|u(t)\|_{H^{\sigma}}\Big{]}\lesssim 1+\int_{0}^{T}\mathbb{E}\Big{[}\|u(t)\|_{L^{\infty}}^{m-2}\|u(t)\|_{H^{\sigma}}\mathds{1}_{t<T^{*}(u_{0})}\Big{]}\mathrm{d}t.$ (4.40)
Proposition 3.5 gives
$\|u-u_{N}\|_{L^{6}([0,\tau],L^{\infty})}+\sup_{t\leq\tau}\|u(t)-u_{N}(t)\|_{H^{\sigma}}\leq CN^{\sigma-\sigma^{\prime}}$ (4.41)
for $\tau<T^{*}(u_{0})$ and $\sigma^{\prime}<\sigma$ with $C=C(u_{0})>0$ a random constant hence Fatou’s lemma gives
$\mathbb{E}\Big{[}\|u(t)\|_{L^{\infty}}^{m-2}\|u(t)\|_{H^{\sigma}}\mathds{1}_{t<T^{*}(u_{0})}\Big{]}\leq\liminf_{N}\mathbb{E}\Big{[}\|u_{N}(t)\|_{L^{\infty}}^{m-2}\|u_{N}(t)\|_{H^{\sigma}}\mathds{1}_{t<T^{*}(u_{0})}\Big{]}$ (4.42)
for almost all $t\in\mathbb{R}$. However, $u_{N}(t)$ has a law $\Pi_{N}\nu$ or $\Pi_{N}\nu_{B}$ independant of $t$ with needed norm with uniformly bounded moment with respect to $N$ hence
$\sup_{N\geq 1}\mathbb{E}\Big{[}\|u_{N}(t)\|_{L^{\infty}}^{m-2}\|u_{N}(t)\|_{H^{\sigma}}\Big{]}<\infty.$ (4.43)
We get
$\mathbb{E}\Big{[}\sup_{t\leq T\wedge T^{*}(u_{0})}\|u(t)\|_{H^{\sigma}}\Big{]}\lesssim 1+T$ (4.44)
which guarantees with the blow up criterion that $T^{*}(u_{0})>T$ almost surely. Since $T>0$ is chosen arbitrary, we get that $T^{*}(u_{0})=+\infty$ and the equation is globally well-posed for $\nu$-almost all initial data. The invariance of the measure follows from the invariance of the truncated measure with the almost sure convergence of $u_{N}$ to $u$ in $C([0,T],H^{\sigma})$ for any $T>0$.
@@
This result is enough to have a well-defined flow defined on a set of full measure. Indeed, it gives a set
$S_{N}=:\{\omega\in\Omega\ ;\ \sup_{-N\leq t\leq N}\|u(\omega,t)\|_{H^{\sigma}}<\infty\}$ (4.45)
of full measure on which solutions are defined up to time $|t|\leq N$, the arguments also work for negative time. Then
$S=\bigcap_{N\geq 1}S_{N}$ (4.46)
yields a set of full measure on which a global flow is well-defined.